OpenStudy (anonymous):

through: (-1,1), parallel to y=4x-1

OpenStudy (anonymous):

y-1=4(x+1)

OpenStudy (anonymous):

could

OpenStudy (anonymous):

you show me how?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

point slope formula y-y1=slope(x-x1)

OpenStudy (anonymous):

slope is the same and just plug in the points from the coordinates given

OpenStudy (anonymous):

got it! thank you :)

OpenStudy (anonymous):

no problem